OpenStudy (skatingdj):

1. By graphing the system of constraints, find the values of x and y that maximize the objective function. 2<=x<=6 1<=y<=5 x+y<=8 maximum for p=3x+2y A. (2,1) B. (6,2) C. (2,5) D. (3,5) 2. By graphing the system of constraints, find the values of x and y that maximize the objective function. 3x+y<=7 x+2y<=9 x=>0 y=>0 Maximum for p=2x+y a. p=3 b. p=4 c. p=6 d. p=10 3.By graphing the system of constraints, find the values of x and y that maximize the objective function. x+y<=11 2y=>x x=>0 y=>0 maximum for p=3x+2y a. p=10 1/3

563blackghost (563blackghost):

Use https://www.calorful.com/graphing/calculator to help you plot the constraints. Once you plug in where they cross is a given point.

OpenStudy (skatingdj):

1. By graphing the system of constraints, find the values of x and y that maximize the objective function. 2<=x<=6 1<=y<=5 x+y<=8 maximum for p=3x+2y A. (2,1) B. (6,2) C. (2,5) D. (3,5) 2. By graphing the system of constraints, find the values of x and y that maximize the objective function. 3x+y<=7 x+2y<=9 x=>0 y=>0 Maximum for p=2x+y a. p=3 b. p=4 c. p=6 d. p=10 3.By graphing the system of constraints, find the values of x and y that maximize the objective function. x+y<=11 2y=>x x=>0 y=>0 maximum for p=3x+2y a. p=10 1/3 b. p=7 1/3 c. p=21 d. p=29 1/3 4.By graphing the system of constraints, find the values of x and y that maximize the objective function. 25<=x<=75 y<=110 8x+6y=>720 y=>0 minimum for c=8x+5y a. c=100 b. c=225 2/3 c. c=633 1/3 d. c=86 2/3 5. By graphing the system of constraints, find the values of x and y that maximize the objective function. 2x+y<=300 x+y<=200 x=>0 y=>0 maximum for p=x+2y a. p=100 b. p=200 c. p=400 d. p=550 These are the last five of a test I'm working on and my brain is getting stuck:/ Any help is appreciated! Thanks

563blackghost (563blackghost):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SkatingDJ These are the last five of a test I'm working on and my brain is getting stuck:/ Any help is appreciated! Thanks \(\color{#0cbb34}{\text{End of Quote}}\) Oooo we can not help with test/exams sorry :(